OpenStudy (anonymous):

x^-2+x^-4= I got x^1/2+x^1/4 but i dont remember the rules well for exponents

OpenStudy (freckles):

x^(-2) doesn't mean x^(1/2) though you can write x^(-2) as 1/x^2

OpenStudy (freckles):

and same for the x^(-4) this is not x^(1/4)

OpenStudy (anonymous):

ohhh okay so do i multiply the exponents?

OpenStudy (jhannybean):

\[x^{-2} -x^{-4}\] You're just looking to factor out this function?

OpenStudy (anonymous):

well for the exponents, it says to put it in A/B

OpenStudy (jhannybean):

Oh okay, so as @freckles mentioned :)

OpenStudy (jhannybean):

Recall that \(x^{-\#} \iff \dfrac{1}{x^{\#}}\) Therefore we can write our function over 1 to make the exponents of the variables positive :) and i made a typo, my function should have read \(x^{-2} + x^{-4}\)*

OpenStudy (anonymous):

i got that now! so do we multiply the exponents to get 1/x^8?

OpenStudy (anonymous):

i completely forgot the rules

OpenStudy (anonymous):

then how?

OpenStudy (jhannybean):

All we can do now, is factor out the denominator.

OpenStudy (freckles):

\[x^{-2}+x^{-4}=\frac{1}{x^2}+\frac{1}{x^4}\] you can combine the fractions by finding a common denominator

OpenStudy (jhannybean):

D'oh!

OpenStudy (jhannybean):

I made a really big error. My mistake was in putting the function as a whole in the denominator. In order to turn each variable positive, you must treat each variable as a separate function.

OpenStudy (jhannybean):

From what @freckles wrote, just find the greatest common denominator between \(x^2\) and \(x^4\).

OpenStudy (anonymous):

?

OpenStudy (anonymous):

common denominator is x^4?

OpenStudy (freckles):

that is right \[x^{-2}+x^{-4} \\ \frac{1}{x^2}+\frac{1}{x^4} \\ \frac{1(x^2)}{x^2(x^2)}+\frac{1}{x^4}\] now you have the same denominator combine the fractions

OpenStudy (anonymous):

so itll be x^2/x^4 do i reduce?

OpenStudy (freckles):

you will have \[\frac{x^2+1}{x^4}\]

OpenStudy (anonymous):

wouldnt that equal to x+1/x^2

OpenStudy (freckles):

no just (x^2+1)/x^4 can't be reduced because you don't have all terms with factor x^2

OpenStudy (anonymous):

than would A/B be 2/4? cause the question asked put the exponents as A/B

OpenStudy (freckles):

what?

OpenStudy (freckles):

are you saying we are suppose to write in x^(A/B) form?

OpenStudy (anonymous):

yes

OpenStudy (freckles):

\[x^{-2}+x^{-4}=x^{\frac{-2}{1}}+x^{\frac{-4}{1}}\] there is both terms written in x^(A/B) form

OpenStudy (anonymous):

oh okay thats easier to understand

OpenStudy (anonymous):

then would it be -6

OpenStudy (freckles):

I don't get your question can you post your whole question

OpenStudy (anonymous):

simplofy the rational expression in the form of A/B. x^-2+x^-4

OpenStudy (freckles):

oh in the form A/B well we did that \[\frac{x^2+1}{x^4} \\ \text{ we have } A=x^2+1 \text{ and } B=x^4\]

OpenStudy (anonymous):

oh my god thank you!

OpenStudy (freckles):

np